Four capacitors are connected in series across a 12.0-V battery. Their capacitances are 1.00, 2.00, 4.00 and 4.00 mF. What must be the capacitance of a fifth capacitor that when added in series to the circuit reduces the voltage across the 1.00 mF capacitor by 1.00 V?
Answer #1

V = Voltage of the battery = 12 volts

Total capacitance is given as

1/Ctotal = 1/C1 + 1/C2 + 1/C3 + 1/C4

1/Ctotal = 1/1 + 1/2 + 1/4 + 1/4

Ctotal = = 0.5 mF

Q1 = charge through C1 = Ctotal V = (0.5 x 10-3) 12 = 6 x 10-3 C

V1 = Voltage across C1 = Q1/C1 = 6 x 10-3 /(1 x 10-3) = 6 volts

V1' = reduced Voltage = 5 volts

Q1' = new charge in the circuit = C1 V1' = (1 x 10-3) 5 = 5 x 10-3 C

V = 12 volts

new total capacitance , Ctotal' = Q1'/V = (5 x 10-3 )/12 = 0.42 x 10-3

new total capacitance is given as

1/Ctotal' = 1/1 + 1/2 + 1/4 + 1/4 + 1/C

1/(0.42 x 10-3) = 1/1 + 1/2 + 1/4 + 1/4 + 1/C

C = 0.00042